OWG Meeting 11-Feb-2015

From GlueXWiki
Jump to: navigation, search

Location and Time

Room: CC F326-327

Time: 1:30pm-2:30pm


(if problems, call phone in conference room: 757-269-6460)

  1. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 120390084.
  2. To join via a Web Browser, go to the page https://bluejeans.com/120390084.
  3. To join via phone, use one of the following numbers and the Conference ID: 120390084
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. More information on connecting to bluejeans is available.

*note: you'll be muted when first connected to the meeting. press *4 to unmute


  1. Announcements
  2. Collaboration Meeting
  3. Selectable "Run Type"s (Paul M.)
  4. Major systems reports
    • Trigger
    • DAQ
    • Monitoring
    • Controls

Previous Meeting


Attendees: David L., Simon T., Paul M., Curtis M., Alex B., Eric P., Sergey F., Sean D., Sergeui P., Alex S., Nathan S., Dave A., Beni Z.

Collaboration Meeting

  • All speakers except Hovanes were present and indicated agreement with their time slot

Selectable "Run Type"s

  • Paul M. presented a request to have some information gathered from shift workers about the type of run they are taking (or attempting to take).
  • This would be written to the RCDB and made available to the Offline to aide in run selection.
  • Some discussion was had on various aspects of this type of information gathering including implications for L1 trigger configuration and readout configuration.
  • Sergey F. will work to implement some additional GUI elements into the DAQ in order to allow users to convey this type information.
  • It was agreed that some elements should default to an invalid choice in order to force the shift taker to make a conscious choice

Major Systems Reports


  • Add SD cards jumpers updated to minimize latency (by ~20 ns)
    • Tagger hall SD cards not yet updated
    • Tested with PS and everything seems to be working
  • Global Trigger test stand in counting house
  • TS busy 2-pin to level convertor not fixed yet (Fernando will do)
  • PS trigger work
  • Scalers use trigger thresholds now instead of readout threshold
  • FCAL: looking at MIP response to help address triggering on noise issue
  • Synch events:
    • Alex has started implementing sync events into the DAQ system
    • Some work is needed on the ROL to properly check that each module has the correct number of events and leave each module in state where it has no more events
    • If an event mismatch occurs, a marker needs to be inserted into the data stream so it can be properly handled in the offline.


  • Testing after firmware using coda 3.01 (not 3.02)
  • Not trivial to switch CODA versions due to initialization scripts
  • Cody still working on f125 firmware
  • 30kHz 100% in mode 8, up to 50 kHz in mode 7
    • these are for period triggers
    • Hope to run at 10kHz robustly (unknown livetime)


- No significant progress to report. Work will resume now that software review is complete


- No report